evaluate the following integral
∫ e^(2x) / e^(2x) – 2 dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{e^{2x}}{e^{2x}-2}dx
e^{2x}-2 = t
2(e^{2x})dx = dt
I = \frac{1}{2}\int \frac{1}{t}dt
I = \frac{1}{2}log(t)+c
I = \frac{1}{2}log(e^{2x}-2)+c